# Electronic modes induced by spin and charge perturbations in Mott and Kondo insulators

### v1

@@ IV.7.2 Electronic modes induced by charge fluctuation in spin-gapped Mott and Kondo insulators
Figure 10: Electronic excitation from the charge-perturbed states of small $\eta/(2N_{\rm u})$ caused by charge fluctuation in the ladder HM for $U/t=8$ and $t_{\perp}/t=4$ [(a), (b)], bilayer HM for $U/t=16$ and $t_{\perp}/t=8$ [(d), (e)], 1D PAM for $U/t=14$ and $t_{\rm K}/t=4$ [(g), (h)], and 2D PAM for $U/t=28$ and $t_{\rm K}/t=8$ [(j), (k)]. (a), (d), (g), (j) $A_{X}(k_{\parallel},0,\omega)t+A_{X}(k_{\parallel},\pi,\omega)t$ [(a)], ${\bar{A}}_{X}({\bm{k}}_{\parallel},0,\omega)t+{\bar{A}}_{X}({\bm{k}}_{\parallel},\pi,\omega)t$ [(d)], $A_{X}(k,\omega)t$ [(g)], and ${\bar{A}}_{X}({\bm{k}},\omega)t$ [(j)] for $|X\rangle$ of the state corresponding to $\prod_{i}{\tilde{n}}^{c}_{{\bm{q}}_{i}}|{\rm GS}\rangle$ at ${\bm{q}}_{i}=({\bm{0}},\pi)$ in (a) and (d); $\prod_{i}({\tilde{n}}^{c}_{{\bm{q}}_{i}}-{\tilde{n}}^{f}_{{\bm{q}}_{i}})|{\rm GS}\rangle$ at ${\bm{q}}_{i}={\bm{0}}$ in (g) and (j) for $\eta/(2N_{\rm u})=0.05$. Gaussian broadening is used with standard deviations of $0.16t$ [(a), (g)] and $0.32t$ [(d), (j)]. (b), (e), (h), (k) Dispersion relations of electronic excitations from ${\tilde{n}}^{c}_{({\bm{0}},\pi)}|{\rm GS}\rangle$ [(b), (e)] and $({\tilde{n}}^{c}_{\bm{0}}-{\tilde{n}}^{f}_{\bm{0}})|{\rm GS}\rangle$ [(h), (k)] in the effective theory. The dispersion relations of the emergent modes are $\omega=-\varepsilon^{\cal R}_{-{\bm{k}}_{\parallel}}+e^{{\cal C}^{0}}_{\bm{0}}$ (dotted brown curves), $\omega=\varepsilon^{\cal A}_{{\bm{k}}_{\parallel}}-e^{{\cal C}^{0}}_{\bm{0}}$ (dotted blue curves), $\omega=-\varepsilon^{\bar{\cal R}}_{-{\bm{k}}_{\parallel}}+e^{{\cal C}^{0}}_{\bm{0}}$ (dotted orange curves), and $\omega=\varepsilon^{\bar{\cal A}}_{{\bm{k}}_{\parallel}}-e^{{\cal C}^{0}}_{\bm{0}}$ (dotted purple curves), which are shown in the relevant momentum regimes for comparison with the numerical results. Here, $e^{{\cal C}^{0}}_{\bm{0}}$ denotes the excitation energy of the charge mode at ${\bm{k}}_{\parallel}={\bm{0}}$ [red arrows in (c), (f), (i), and (l)]. The dispersion relations of the other modes and continua in (b), (e), (h), and (k) are the same as those in Figs. 3(g), 3(j), 3(h), and 3(k), respectively. The up and down arrows on the right sides of (b), (e), (h), and (k) indicate $\omega$-shifts from the unperturbed modes by $e^{{\cal C}^{0}}_{\bm{0}}$ and $e^{{\cal C}^{0}}_{\bm{0}}$, respectively. (c), (f), (i), (l) Dispersion relations of charge excitations in the effective theory. The model parameters in (c), (f), (i), and (l) are the same as those in (b), (e), (h), and (k), respectively. The red arrows indicate the excitation energy of the charge mode at ${\bm{k}}_{\parallel}={\bm{0}}$, i.e., $e^{{\cal C}^{0}}_{\bm{0}}$. The dispersion relations of the modes and continua in (c), (f), (i), and (l) are the same as those in Figs. 5(b), 5(c), 5(e), and 5(f), respectively. The solid green lines indicate $\omega=0$.
@@ V.1 Outline
In Secs. V.2V.5, the origins, underlying mechanisms, and properties of the magnetization-induced states are elucidated. Section V.2 explains the magnetization-induced states from the viewpoint of quantum numbers. Their emergence is demonstrated in spin-gapped Mott and Kondo insulators using the effective theory for weak inter-unit-cell hopping and numerical calculations in Sec. V.3. Section V.4 shows that electronic modes are not induced inside the band bap by magnetizing insulators with gapless spin excitation. Section V.5 shows that the spectral functions of the magnetized states can also be obtained from those of the corresponding doped states in the models with U-U via the Shiba transformation.
@@ V.3.1 Dispersion relations of magnetization-induced modes in spin-gapped Mott and Kondo insulators
Figure 11: Electronic excitation from the lowest-energy spin-perturbed states of small $S/S_{\rm max}$ in the ladder HM for $U/t=8$ and $t_{\perp}/t=4$ [(a)–(c), (g)–(i)], bilayer HM for $U/t=16$ and $t_{\perp}/t=8$ [(d)–(f), (j)–(l)], 1D PAM for $U/t=14$ and $t_{\rm K}/t=4$ [(m)–(o), (s)–(u)], and 2D PAM for $U/t=28$ and $t_{\rm K}/t=8$ [(p)–(r), (v)–(x)]. (a), (b), (d), (e), (m), (n), (p), (q) $A^{\sigma}_{X}(k_{\parallel},0,\omega)t+A^{\sigma}_{X}(k_{\parallel},\pi,\omega)t$ [(a), (b)], ${\bar{A}}^{\sigma}_{X}({\bm{k}}_{\parallel},0,\omega)t+{\bar{A}}^{\sigma}_{X}({\bm{k}}_{\parallel},\pi,\omega)t$ [(d), (e)], $A^{\sigma}_{X}(k,\omega)t$ [(m), (n)], and ${\bar{A}}^{\sigma}_{X}({\bm{k}},\omega)t$ [(p), (q)] for $|X\rangle$ of the state corresponding to the magnetized ground state of $S/S_{\rm max}=0.05$ and $S^{z}=S$ for $\sigma=\downarrow$ [(a), (d), (m), (p)] and $\uparrow$ [(b), (e), (n), (q)]. (c), (f), (o), (r) $A_{X}(k_{\parallel},0,\omega)t+A_{X}(k_{\parallel},\pi,\omega)t$ [(c)], ${\bar{A}}_{X}({\bm{k}}_{\parallel},0,\omega)t+{\bar{A}}_{X}({\bm{k}}_{\parallel},\pi,\omega)t$ [(f)], $A_{X}(k,\omega)t$ [(o)], and ${\bar{A}}_{X}({\bm{k}},\omega)t$ [(r)] for $|X\rangle$ of the state corresponding to the lowest-energy state of $S/S_{\rm max}=0.05$ and $S^{z}=0$. Gaussian broadening is used with standard deviations of $0.16t$ [(a)–(c), (m)–(o)] and $0.32t$ [(d)–(f), (p)–(r)]. (g)–(l), (s)–(x) Dispersion relations from the ground state of $S/S_{\rm max}\rightarrow+0$ for $\downarrow$-spin electrons [(g), (j), (s), (v)] and $\uparrow$-spin electrons [(h), (k), (t), (w)] under the magnetic field of $H=\Delta_{\rm spin}$ ($S^{z}/S_{\rm max}\rightarrow+0$) [(g), (h), (j), (k), (s), (t), (v), (w)] and without a magnetic field ($S^{z}=0$) [(i), (l), (u), (x)] in the effective theory. The dispersion relations of the emergent modes are $\omega=\varepsilon^{{\cal A}_{\uparrow}}_{{\bm{k}}_{\parallel}+{\bm{\pi}}}-e^{{\cal T}^{1}}_{\bm{\pi}}$ (dotted blue curves) and $\omega=\varepsilon^{\bar{\cal A}_{\uparrow}}_{{\bm{k}}_{\parallel}+{\bm{\pi}}}-e^{{\cal T}^{1}}_{\bm{\pi}}$ (dotted purple curves) in (g), (i), (j), (l), (s), (u), (v), and (x); $\omega=-\varepsilon^{{\cal R}_{\uparrow}}_{-{\bm{k}}_{\parallel}+{\bm{\pi}}}+e^{{\cal T}^{1}}_{\bm{\pi}}$ (dotted brown curves) and $\omega=-\varepsilon^{\bar{\cal R}_{\uparrow}}_{-{\bm{k}}_{\parallel}+{\bm{\pi}}}+e^{{\cal T}^{1}}_{\bm{\pi}}$ (dotted orange curves) in (h), (i), (k), (l), (t), (u), (w), and (x), which are shown in the relevant momentum regimes for comparison with the numerical results. The dispersion relations of the other modes and continua in (g)–(i), (j)–(l), (s)–(u), and (v)–(x) are the same as those in Figs. 3(g), 3(j), 3(h), and 3(k), respectively, with the effective chemical potential of $\mu_{\rm eff}=-\frac{\Delta_{\rm spin}}{2}$ for $\downarrow$-spin electrons [(g), (j), (s), (v)] and $\frac{\Delta_{\rm spin}}{2}$ for $\uparrow$-spin electrons [(h), (k), (t), (w)] under the magnetic field of $H=\Delta_{\rm spin}$ and $\mu_{\rm eff}=0$ [(i), (l), (u), (x)] without a magnetic field. The solid green lines indicate $\omega=0$.
@@
Figure 12: Electronic excitation from the lowest-energy spin-perturbed states of small $S/S_{\rm max}$ in the 1D KLM for $J_{\rm K}/t=4$ [(a)–(c), (g)–(i)] and 2D KLM for $J_{\rm K}/t=8$ [(d)–(f), (j)–(l)]. (a), (b), (d), (e) $A^{\sigma}_{X}({\bm{k}},\omega)t$ for $|X\rangle$ of the state corresponding to the magnetized ground state of $S/S_{\rm max}=0.05$ and $S^{z}=S$ for $\sigma=\downarrow$ [(a), (d)] and $\uparrow$ [(b), (e)]. (c), (f) $A_{X}({\bm{k}},\omega)t$ for $|X\rangle$ of the state corresponding to the lowest-energy state of $S/S_{\rm max}=0.05$ and $S^{z}=0$. Gaussian broadening is used with standard deviations of $0.12t$ [(a)–(c)] and $0.24t$ [(d)–(f)]. (g)–(l) Dispersion relations from the ground state of $S/S_{\rm max}\rightarrow+0$ for $\downarrow$-spin electrons [(g), (j)] and $\uparrow$-spin electrons [(h), (k)] under the magnetic field of $H=\Delta_{\rm spin}$ ($S^{z}/S_{\rm max}\rightarrow+0$) [(g), (h), (j), (k)] and without a magnetic field ($S^{z}=0$) [(i), (l)] in the effective theory. The dispersion relations of the emergent modes are $\omega=\varepsilon^{{\cal A}_{\uparrow}}_{{\bm{k}}_{\parallel}+{\bm{\pi}}}-e^{{\cal T}^{1}}_{\bm{\pi}}$ (dotted blue curves) [(g), (i), (j), (l)] and $\omega=-\varepsilon^{{\cal R}_{\uparrow}}_{-{\bm{k}}_{\parallel}+{\bm{\pi}}}+e^{{\cal T}^{1}}_{\bm{\pi}}$ (dotted brown curves) [(h), (i), (k), (l)], which are shown in the relevant momentum regimes for comparison with the numerical results. The dispersion relations of the other modes and continua in (g)–(i) and (j)–(l) are the same as those in Figs. 3(i) and 3(l), respectively, with the effective chemical potential of $\mu_{\rm eff}=-\frac{\Delta_{\rm spin}}{2}$ for $\downarrow$-spin electrons [(g), (j)] and $\frac{\Delta_{\rm spin}}{2}$ for $\uparrow$-spin electrons [(h), (k)] under the magnetic field of $H=\Delta_{\rm spin}$ and $\mu_{\rm eff}=0$ [(i), (l)] without a magnetic field. The solid green lines indicate $\omega=0$.
@@
To demonstrate the emergence of electronic modes caused by magnetizing spin-gapped Mott and Kondo insulators, we consider the lowest-energy states of $S^{z}=\pm 1$ at half filling in the ladder and bilayer HMs and 1D and 2D PAMs and KLMs as perturbed states, which are obtained as $|{\cal T}^{\pm 1}\rangle_{{\bm{q}}_{\parallel}}$ [Eq. (27)] in the effective theory. In these models, ${\bm{q}}_{\parallel}={\bm{\pi}}$, and the spin gap is obtained as $\Delta_{\rm spin}=e^{\cal T}_{\bm{\pi}}$ at $H=0$.
@@
$\omega=-\varepsilon^{{\tilde{\cal R}}_{\sigma}}_{-{\bm{k}}_{\parallel}+{\bm{q}}_{\parallel}}+e^{{\cal T}^{2s^{z}}}_{{\bm{q}}_{\parallel}}\quad\text{in}\quad A^{\sigma}_{{\cal T}^{2s^{z}}_{{\bm{q}}_{\parallel}}}({\bm{k}},\omega),$ (81)
@@
In the ladder and bilayer HMs and 1D and 2D PAMs and KLMs, electronic modes originating from the $\cal A$$_{\uparrow}$ mode emerge within the band gap in the electron-addition spectrum for $\downarrow$-spin electrons, exhibiting the dispersion relation of Eq. (80) for ${\tilde{\cal A}}_{\sigma}={\cal A}_{\uparrow}$ (Tables 14; Appendix B) [dotted bule curves in Figs. 11(g), 11(j), 11(s), 11(v), 12(g), and 12(j)] from the magnetized ground state of $S^{z}=1$ ($S^{z}/S_{\rm max}\rightarrow+0$), which are consistent with the numerical results [Figs. 11(a), 11(d), 11(m), 11(p), 12(a), and 12(d)].
@@
Even though the magnetization-induced modes emerge under the magnetic field, the system remains insulating in spin-gapped Mott and Kondo insulators at $\mu=0$. The chemical potential required for insulator–metal transition under the magnetic field does not change between the cases with and without the magnetization-induced modes. These characteristics are explained as follows.
@@
V.4 Absence of magnetization-induced ingap modes in spin-gapless Mott and Kondo insulators
Although electronic modes are induced within the band bap by magnetizing spin-gapped Mott and Kondo insulators, electronic modes are not induced inside the band gap by magnetizing insulators with gapless spin excitation.

### v2

@@ IV.7.2 Electronic modes induced by charge fluctuation in spin-gapped Mott and Kondo insulators
Figure 10: Electronic excitation from the charge-perturbed states of small $\eta/(2N_{\rm u})$ caused by charge fluctuation in the ladder HM [(a), (b)], bilayer HM [(d), (e)], 1D PAM [(g), (h)], and 2D PAM [(j), (k)]. (a), (d), (g), (j) $A_{X}(k_{\parallel},0,\omega)t+A_{X}(k_{\parallel},\pi,\omega)t$ [(a)], ${\bar{A}}_{X}({\bm{k}}_{\parallel},0,\omega)t+{\bar{A}}_{X}({\bm{k}}_{\parallel},\pi,\omega)t$ [(d)], $A_{X}(k,\omega)t$ [(g)], and ${\bar{A}}_{X}({\bm{k}},\omega)t$ [(j)] for $|X\rangle$ of the state corresponding to $\prod_{i}{\tilde{n}}^{c}_{{\bm{q}}_{i}}|{\rm GS}\rangle$ at ${\bm{q}}_{i}=({\bm{0}},\pi)$ in (a) and (d); $\prod_{i}({\tilde{n}}^{c}_{{\bm{q}}_{i}}-{\tilde{n}}^{f}_{{\bm{q}}_{i}})|{\rm GS}\rangle$ at ${\bm{q}}_{i}={\bm{0}}$ in (g) and (j) for $\eta/(2N_{\rm u})=0.05$. (b), (e), (h), (k) Dispersion relations of electronic excitations from ${\tilde{n}}^{c}_{({\bm{0}},\pi)}|{\rm GS}\rangle$ [(b), (e)] and $({\tilde{n}}^{c}_{\bm{0}}-{\tilde{n}}^{f}_{\bm{0}})|{\rm GS}\rangle$ [(h), (k)] in the effective theory. The dispersion relations of the emergent modes are $\omega=-\varepsilon^{\cal R}_{-{\bm{k}}_{\parallel}}+e^{{\cal C}^{0}}_{\bm{0}}$ (dotted brown curves), $\omega=\varepsilon^{\cal A}_{{\bm{k}}_{\parallel}}-e^{{\cal C}^{0}}_{\bm{0}}$ (dotted blue curves), $\omega=-\varepsilon^{\bar{\cal R}}_{-{\bm{k}}_{\parallel}}+e^{{\cal C}^{0}}_{\bm{0}}$ (dotted orange curves), and $\omega=\varepsilon^{\bar{\cal A}}_{{\bm{k}}_{\parallel}}-e^{{\cal C}^{0}}_{\bm{0}}$ (dotted purple curves), which are shown in the relevant momentum regimes for comparison with the numerical results. Here, $e^{{\cal C}^{0}}_{\bm{0}}$ denotes the excitation energy of the charge mode at ${\bm{k}}_{\parallel}={\bm{0}}$ [red arrows in (c), (f), (i), and (l)]. The dispersion relations of the other modes and continua in (b), (e), (h), and (k) are the same as those in Figs. 3(g), 3(j), 3(h), and 3(k), respectively. The up and down arrows on the right sides of (b), (e), (h), and (k) indicate $\omega$-shifts from the unperturbed modes by $e^{{\cal C}^{0}}_{\bm{0}}$ and $e^{{\cal C}^{0}}_{\bm{0}}$, respectively. (c), (f), (i), (l) Dispersion relations of charge excitations in the effective theory in the ladder HM [(c)], bilayer HM [(f)], 1D PAM [(i)], and 2D PAM [(l)]. The red arrows indicate the excitation energy of the charge mode at ${\bm{k}}_{\parallel}={\bm{0}}$, i.e., $e^{{\cal C}^{0}}_{\bm{0}}$. The dispersion relations of the modes and continua in (c), (f), (i), and (l) are the same as those in Figs. 5(b), 5(c), 5(e), and 5(f), respectively. The solid green lines indicate $\omega=0$.
@@ V.1 Outline
In Secs. V.2V.5, the origins, underlying mechanisms, and properties of the magnetization-induced states are elucidated. Section V.2 explains the magnetization-induced states from the viewpoint of quantum numbers. Their emergence is demonstrated in spin-gapped Mott and Kondo insulators using the effective theory for weak inter-unit-cell hopping and numerical calculations in Sec. V.3. Section V.4 shows that electronic modes are not induced inside the band gap by magnetizing insulators with gapless spin excitation. Section V.5 shows that the spectral functions of the magnetized states can also be obtained from those of the corresponding doped states in the models with U-U via the Shiba transformation.
@@ V.3.1 Dispersion relations of magnetization-induced modes in spin-gapped Mott and Kondo insulators
Figure 11: Electronic excitation from the lowest-energy spin-perturbed states of small $S/S_{\rm max}$ in the ladder HM [(a)–(c), (g)–(i)], bilayer HM [(d)–(f), (j)–(l)], 1D PAM [(m)–(o), (s)–(u)], and 2D PAM [(p)–(r), (v)–(x)]. (a), (b), (d), (e), (m), (n), (p), (q) $A^{\sigma}_{X}(k_{\parallel},0,\omega)t+A^{\sigma}_{X}(k_{\parallel},\pi,\omega)t$ [(a), (b)], ${\bar{A}}^{\sigma}_{X}({\bm{k}}_{\parallel},0,\omega)t+{\bar{A}}^{\sigma}_{X}({\bm{k}}_{\parallel},\pi,\omega)t$ [(d), (e)], $A^{\sigma}_{X}(k,\omega)t$ [(m), (n)], and ${\bar{A}}^{\sigma}_{X}({\bm{k}},\omega)t$ [(p), (q)] for $|X\rangle$ of the state corresponding to the magnetized ground state of $S/S_{\rm max}=0.05$ and $S^{z}=S$ for $\sigma=\downarrow$ [(a), (d), (m), (p)] and $\uparrow$ [(b), (e), (n), (q)]. (c), (f), (o), (r) $A_{X}(k_{\parallel},0,\omega)t+A_{X}(k_{\parallel},\pi,\omega)t$ [(c)], ${\bar{A}}_{X}({\bm{k}}_{\parallel},0,\omega)t+{\bar{A}}_{X}({\bm{k}}_{\parallel},\pi,\omega)t$ [(f)], $A_{X}(k,\omega)t$ [(o)], and ${\bar{A}}_{X}({\bm{k}},\omega)t$ [(r)] for $|X\rangle$ of the state corresponding to the lowest-energy state of $S/S_{\rm max}=0.05$ and $S^{z}=0$. (g)–(l), (s)–(x) Dispersion relations from the ground state of $S/S_{\rm max}\rightarrow+0$ for $\downarrow$-spin electrons [(g), (j), (s), (v)] and $\uparrow$-spin electrons [(h), (k), (t), (w)] under the magnetic field of $H=\Delta_{\rm spin}$ ($S^{z}/S_{\rm max}\rightarrow+0$) [(g), (h), (j), (k), (s), (t), (v), (w)] and without a magnetic field ($S^{z}=0$) [(i), (l), (u), (x)] in the effective theory. The dispersion relations of the emergent modes are $\omega=\varepsilon^{{\cal A}_{\uparrow}}_{{\bm{k}}_{\parallel}+{\bm{\pi}}}-e^{{\cal T}^{1}}_{\bm{\pi}}$ (dotted blue curves) and $\omega=\varepsilon^{\bar{\cal A}_{\uparrow}}_{{\bm{k}}_{\parallel}+{\bm{\pi}}}-e^{{\cal T}^{1}}_{\bm{\pi}}$ (dotted purple curves) in (g), (i), (j), (l), (s), (u), (v), and (x); $\omega=-\varepsilon^{{\cal R}_{\uparrow}}_{-{\bm{k}}_{\parallel}+{\bm{\pi}}}+e^{{\cal T}^{1}}_{\bm{\pi}}$ (dotted brown curves) and $\omega=-\varepsilon^{\bar{\cal R}_{\uparrow}}_{-{\bm{k}}_{\parallel}+{\bm{\pi}}}+e^{{\cal T}^{1}}_{\bm{\pi}}$ (dotted orange curves) in (h), (i), (k), (l), (t), (u), (w), and (x), which are shown in the relevant momentum regimes for comparison with the numerical results. The dispersion relations of the other modes and continua in (g)–(i), (j)–(l), (s)–(u), and (v)–(x) are the same as those in Figs. 3(g), 3(j), 3(h), and 3(k), respectively, with the effective chemical potential of $\mu_{\rm eff}=-\frac{\Delta_{\rm spin}}{2}$ for $\downarrow$-spin electrons [(g), (j), (s), (v)] and $\frac{\Delta_{\rm spin}}{2}$ for $\uparrow$-spin electrons [(h), (k), (t), (w)] under the magnetic field of $H=\Delta_{\rm spin}$ and $\mu_{\rm eff}=0$ [(i), (l), (u), (x)] without a magnetic field. The solid green lines indicate $\omega=0$.
@@
Figure 12: Electronic excitation from the lowest-energy spin-perturbed states of small $S/S_{\rm max}$ in the 1D KLM [(a)–(c), (g)–(i)] and 2D KLM [(d)–(f), (j)–(l)]. (a), (b), (d), (e) $A^{\sigma}_{X}({\bm{k}},\omega)t$ for $|X\rangle$ of the state corresponding to the magnetized ground state of $S/S_{\rm max}=0.05$ and $S^{z}=S$ for $\sigma=\downarrow$ [(a), (d)] and $\uparrow$ [(b), (e)]. (c), (f) $A_{X}({\bm{k}},\omega)t$ for $|X\rangle$ of the state corresponding to the lowest-energy state of $S/S_{\rm max}=0.05$ and $S^{z}=0$. (g)–(l) Dispersion relations from the ground state of $S/S_{\rm max}\rightarrow+0$ for $\downarrow$-spin electrons [(g), (j)] and $\uparrow$-spin electrons [(h), (k)] under the magnetic field of $H=\Delta_{\rm spin}$ ($S^{z}/S_{\rm max}\rightarrow+0$) [(g), (h), (j), (k)] and without a magnetic field ($S^{z}=0$) [(i), (l)] in the effective theory. The dispersion relations of the emergent modes are $\omega=\varepsilon^{{\cal A}_{\uparrow}}_{{\bm{k}}_{\parallel}+{\bm{\pi}}}-e^{{\cal T}^{1}}_{\bm{\pi}}$ (dotted blue curves) [(g), (i), (j), (l)] and $\omega=-\varepsilon^{{\cal R}_{\uparrow}}_{-{\bm{k}}_{\parallel}+{\bm{\pi}}}+e^{{\cal T}^{1}}_{\bm{\pi}}$ (dotted brown curves) [(h), (i), (k), (l)], which are shown in the relevant momentum regimes for comparison with the numerical results. The dispersion relations of the other modes and continua in (g)–(i) and (j)–(l) are the same as those in Figs. 3(i) and 3(l), respectively, with the effective chemical potential of $\mu_{\rm eff}=-\frac{\Delta_{\rm spin}}{2}$ for $\downarrow$-spin electrons [(g), (j)] and $\frac{\Delta_{\rm spin}}{2}$ for $\uparrow$-spin electrons [(h), (k)] under the magnetic field of $H=\Delta_{\rm spin}$ and $\mu_{\rm eff}=0$ [(i), (l)] without a magnetic field. The solid green lines indicate $\omega=0$.
@@
To demonstrate the emergence of electronic modes caused by magnetizing spin-gapped Mott and Kondo insulators, we consider the lowest-energy states of $S^{z}=\pm 1$ at half filling in the ladder and bilayer HMs and 1D and 2D PAMs and KLMs as perturbed states, which are obtained as $|{\cal T}^{\pm 1}\rangle_{{\bm{q}}_{\parallel}}$ [Eq. (29)] in the effective theory. In these models, ${\bm{q}}_{\parallel}={\bm{\pi}}$, and the spin gap is obtained as $\Delta_{\rm spin}=e^{\cal T}_{\bm{\pi}}$ at $H=0$ (Fig. 4).
@@
$\omega=-\varepsilon^{{\tilde{\cal R}}_{\sigma}}_{-{\bm{k}}_{\parallel}+{\bm{q}}_{\parallel}}+e^{{\cal T}^{2s^{z}}}_{{\bm{q}}_{\parallel}}\quad\text{in}\quad A^{\sigma}_{{\cal T}^{2s^{z}}_{{\bm{q}}_{\parallel}}}({\bm{k}},\omega)$ (83)
@@
In the ladder and bilayer HMs and 1D and 2D PAMs and KLMs, electronic modes originating from the $\cal A$$_{\uparrow}$ mode emerge within the band gap in the electron-addition spectrum for $\downarrow$-spin electrons, exhibiting the dispersion relation of Eq. (82) for ${\tilde{\cal A}}_{\sigma}={\cal A}_{\uparrow}$ (Tables 14; Appendix B) [dotted blue curves in Figs. 11(g), 11(j), 11(s), 11(v), 12(g), and 12(j)] from the magnetized ground state of $S^{z}=1$ ($S^{z}/S_{\rm max}\rightarrow+0$), which are consistent with the numerical results [Figs. 11(a), 11(d), 11(m), 11(p), 12(a), and 12(d)].
@@
Even though the magnetization-induced modes emerge under the magnetic field, the system remains insulating in spin-gapped Mott and Kondo insulators at $\mu=0$. The chemical potential required for insulator–metal transition under the magnetic field does not change between the cases with and without the magnetization-induced modes. These characteristics are explained as follows:
@@
V.4 Absence of magnetization-induced in-gap modes in spin-gapless Mott and Kondo insulators
Although electronic modes are induced within the band gap by magnetizing spin-gapped Mott and Kondo insulators, electronic modes are not induced inside the band gap by magnetizing insulators with gapless spin excitation.